# Chemical Sciences Division, Lawrence Berkeley National Laboratory, Berkeley, CA 94720, USA

### v1

@@ 4.3 Property-Specific Benchmark Analysis
For thermochemistry (TC) (Figure 2), standard MAEs typically fall between 1 and 5 kcal/mol, indicating that even the best hybrid functionals still fail to reach chemical accuracy (1 kcal/mol) for most sets. Particularly large errors are observed for AE11 (atomic energies of heavy elements), MX34 (atomization energies of NaCl-like ionic clusters), and MB16-43 (mindless benchmarks). Ionization potential (IP) data sets, including DIPCS9 (double IP), generally show MAEs around 2–3 kcal/mol, while electron affinity (EA) sets average near 2 kcal/mol. Among all tested functionals, CF22D and B97M-V perform best, with CF22D showing clear advantages on challenging sets like MB16-43 and TAE_W4-17MR. Notably, although B97M-V and B97X-V perform similarly on IP and EA, B97X-V outperforms B97M-V on time-dependent DFT (TDDFT) excitation energies 209, suggesting little correlation between accuracy in adiabatic IP/EA and accuracy for TDDFT excitation energies (which depends on the additional adiabatic approximation). It is possible that orbital-optimized DFT (OO-DFT) excitation energies210 might show stronger correlations. Furthermore, B97M-V and B97X-V yield reasonable performance on MX34 but greatly overestimate solid-state band gaps 211, implying that MX34 may not fully capture periodic system behavior (we also note that solid state gaps are eigenvalue differences while MX34 are total energy differences).
@@
Barrier‑height (BH) data (Figure 5) tell a more comforting story: most hybrids sit between 1 and 2 kcal mol-1. That might look like an improvement over TC, but remember that BH reference values are usually smaller in this database, so the same absolute error can hide larger percentage mistakes. B97M‑V and CF22D perform best in this category, though CF22D shows large errors on the cycloreversion subset CRBH14.
@@
For transition metal (TM) systems (Figure 5), top-performing functionals achieve MAEs around 1–3 kcal/mol for organometallic complexes, but significantly higher errors (5 kcal/mol) occur for metal-dominated sets such as 3d4dIPSS, CUAGAU83, and TMD10. Performance varies with reaction type. CF22D performs poorly on reaction energies in MOR13 and ROST61, despite having the lowest overall NER for TM. In contrast, B97M-V excels for organometallic reaction energies and barriers but struggles with small cluster sets like CUAGAU83 and DAPD. Interestingly, the performance gap between empirical and non-empirical functionals narrows for TM systems, suggesting reduced advantages of empirical parameterization, at least when extrapolating beyond training data.
@@
For noncovalent interactions (NC) (Figure 3), sub-kcal/mol accuracy is generally attainable except for radical systems (TA13, O24). B97M-V, B97X-V, and B97M-V lead their respective rungs, echoing findings from MGCDB84 and GMTKN55. Notably, they show low errors for ionic hydration (e.g., FmH2O10, HW6Cl, HW6F), possibly reflecting overfitting. Conversely, they perform poorly on the 3B-69 set, indicating that VV10 lacks the explicit three-body terms captured by DFT-D4. When three-body terms from B97M-D4 are added, the NERs of B97M-V improve from 1.78 to 1.21 (3B-69) and 1.11 to 0.98 (3BHET) though the overall mean NER on NC is not improved. We therefore recommend adding three‑body dispersion to VV10‑based DFAs for large systems.
Isomerization (ISO) sets (Figure 4) behave much like NC: small molecules are easy, but errors appear to grow with size and topological complexity. B97M‑V is still the best overall, particularly for potential‑energy curves such as A19Rel6 and S66Rel7. However, it fails badly on C60ISO8, consistent with known problems for range-separated hybrids on fullerene isomer energies.212, 213 Interestingly, B97M(2) remedies this issue, hinting that cancellation between exchange and correlation in medium-to-long range is essential. CF22D ranks second overall; B97X-V ranks third but shows inconsistent results, performing well on ACONF (alkane conformers) but poorly on AlkIsomer11 (constitutional alkane isomers).
@@
For vibrational frequencies (Figure 5), the typical MAE is around 30 , with the best-performing functional, PBE0‑D4, reducing this to 21.8 . For dipole moments, the mean absolute regularized error is about 3.5%. Static polarizabilities show a standard MARE of 2.4% on Pol130 but only about 1.3% on HR46 and T144, reflecting the greater challenge of spin-polarized or exotic molecules in Pol130. For orientation-dependent electric fields (OEEF), the standard MARE rises to 8.8%, indicating that stronger fields amplify functional errors. Interestingly, for electric field properties (EF) (Figure 5), B97M-V and CF22D clearly underperform others, contrary to trends in ground-state energy performance. This reversal reinforces the concern that B97M-V and CF22D may overfit energy-based training sets and highlights the need to include molecular response properties during functional development to ensure broader transferability.
@@ 4.4 Additional Considerations for Double Hybrid Functionals
Compared to hybrid functionals, double-hybrid (DH) functionals offer substantial improvements across most data sets. The overall mean NERs of B97M(2) and DSD-PBE-D4 are 0.79 and 1.29, respectively—representing relative improvements of 27% and 43% compared to their hybrid analogues, B97M-V and PBE0-D4. However, these gains come with new challenges due to the inclusion of MP2 correlation, which introduces sensitivity to computational choices such as frozen-core treatment, basis set size, and multi-reference (MR) character.
One important consideration is the use of frozen core (FC) approximations. Since the semi-local component of a DH is fitted with a specific FC protocol, calculations should, in principle, replicate that protocol. For example, B97M(2) was trained with the FC approximation. If this is omitted, the MAE on the AlkAtom19 set rises dramatically from 0.18 kcal/mol to 6.24 kcal/mol—an increase by a factor of 35. However, the problem becomes more complicated when considering elements outside the training set or software-specific definitions of frozen orbitals. For instance, in Q-Chem, the K+ ion is treated with 9 frozen orbitals (a very large core!), leading to an 8.15 kcal/mol MAE on the CHB6 set. Without FC, the error drops to just 1.23 kcal/mol. Another extreme case is AE11, where enabling FC inflates the error by more than two orders of magnitude. To avoid such artifacts, we report all-electron results for B97M(2) whenever the original FC recipe is demonstrably inadequate (e.g. HEAVY28_13, HEAVY28_14, CHB6_3, CHB6_6, MB16-43, AE11, MX34, and all TM-related sets).
DSD-PBE-D4 was originally developed with a different FC protocol: valence electrons are kept, and outer-core and orbitals are included for transition metals and heavier alkali/alkaline earth atoms. This scheme is difficult to automate in our setup. However, DSD-PBE-D4 is a non-empirical functional, and its performance appears less sensitive to FC choices. Therefore, we choose not to apply any FC approximation for DSD-PBE-D4 in this study. Additionally, as the MP2 part in modern DH functionals (especially RI-MP2 or localized RI-MP2) is often faster than the SCF stage, the FC technique might not be necessary for practical speedups and should perhaps be avoided altogether in both training and deployment (although we must admit this might then require basis sets capable of describing core-valence and core correlation).
@@
Another important source of error in DH-DFT is basis-set incompleteness. The MP2 component converges only as with the maximum angular momentum , whereas the SCF part approaches the CBS limit exponentially. 214 A quadruple-zeta basis is therefore still far from complete for a double hybrid. In practice one should use the same basis that was employed during parametrization. Accordingly, we pair B97M(2) with def2-QZVPPD whenever possible. An instructive exception is the AE11 atomic set: a specially designed core-correlated basis reduces the error much more than def2-QZVPPD, indicating that explicit CBS extrapolation or custom basis optimization can be essential for the most demanding cases. The result also suggests that future double hybrids should be trained with larger, explicitly CBS-compatible basis sets, if possible.
Multi-reference (MR) effects pose an additional challenge. DSD-PBE-D4 outperforms PBE0-D4 on most single-reference sets, yet it degrades on strongly MR systems such as the TAE_W4-17MR and ORBH36 subsets. While no practical density functional can yet quantitatively describe strong correlations, semilocal DFT can describe some non-dynamical correlation because its “exchange” component mimics MR behavior,215 whereas the perturbative MP2 term is highly sensitive to any MR character. Consequently, double hybrids react strongly to small changes in the reference determinant. The HNS polarizability illustrates the point: when the electric field lies in the molecular plane, rises by 0.04 and the DSD-PBE-D4 error becomes enormous (Pol132_153 and Pol132_154); when the field is perpendicular, changes by only 0.008 and the error drops by a factor of twelve (Pol132_155), though it remains larger than for hybrids. B97M(2) is less vulnerable to this problem, but its improvement over B97M-V on MR-dominated data is still smaller than on single-reference cases. Beyond-MP2 correlation, such as the random phase approximation (RPA)216 or regularized MP2217, 218 could potentially reduce this extreme sensitivity.
@@ 5 Conclusions
We have presented GSCDB138, a rigorously curated suite of 138 benchmark sets that roughly doubles the data-point count of GMTKN55/MGCDB84 (to 8,383 energies) while broadening the chemical and physical scope. New components include realistic transition-metal thermochemistry and kinetics, three-body non-covalent interactions, dipoles, polarizabilities, oriented-field response energies, and dimer vibrational frequencies. At the same time, legacy data have been re-evaluated, obsolete reference values replaced, and spin-contaminated or redundant entries removed. By gathering all of these elements under a single, fully documented umbrella, GSCDB138 furnishes a coherent, high-accuracy platform for functional assessment today and a robust training ground for the next generation of non-empirical or machine-learned density functionals.
Benchmarking 29 broadly used density functionals on this database confirms the expected Jacob’s-ladder hierarchy but also highlights important exceptions. Double hybrids remain the most accurate rung, with B97M(2) and DSD-PBE-D4 reducing the average error by roughly one-quarter relative to the best hybrids. Among hybrids, B97M-V emerges as the most transferable HMGGA, while B97X-V is the most balanced HGGA, especially for electric-field and vibrational properties. B97M-V is the clear leader among meta-GGAs, and revPBE-D4 is the safest choice within the conventional GGA class. However, some categories violate the ladder ordering: semi-local r2SCAN-D4 outperforms many hybrids on frequencies, and EF errors show little correlation with ground-state energetics. These findings warn against relying on energy-only training sets and underscore the need to include density-sensitive and response quantities in future functional design. From an applications standpoint, the present study suggests the following guidelines. For main-group chemistry or organometallic reactions, B97M-V (or its D4 variant, or the runner-up, CF22D, where VV10 is unavailable) offers the best accuracy. CF22D is also a valuable choice for metal clusters. For frequencies, B3LYP-D4, PBE0-D4, and r2SCAN-D4 are recommended.
Double-hybrid accuracy is impressive but not turnkey. We show that frozen-core protocols, basis-set completeness, and potential multi-reference character can each change double-hybrid errors by a large ratio (even orders of magnitude). In practical calculations we therefore recommend (i) reproducing the training frozen-core definition whenever possible, (ii) using at least the training basis—def2-QZVPPD for B97M(2)—and (iii) monitoring and other MR diagnostics for reactions that might challenge perturbative correlation. When such precautions are unfeasible, a robust hybrid (e.g., B97X-V or B97M-V) may be the safer alternative at present.
@@
This work was supported by the Director, Office of Science, Office of Basic Energy Sciences, of the U.S. Department of Energy through the Gas Phase Chemical Physics Program, under Contract No. DE-AC02-05CH11231. We thank the authors of MGCDB84, GMTKN55, and ACCDB for providing a solid foundation for this work. We are deeply grateful to Diptarka Hait for insightful discussions and valuable suggestions on refining the dipole moment and static polarizability data sets. We also thank Bun Chan for his generous guidance on the use of his benchmark sets and for providing so much highly accurate reference data. Our appreciation extends to Tarek Scheele and Tim Neudecker for clarifying details of the OEEF set.
Most importantly, we thank all researchers who devote their time and effort to generating high-quality reference values, whether or not their work is included in this database. We fully recognize not only the substantial computational resources to complete high-level wavefunction calculations, as well as the human time required in the meticulous and often tedious process of evaluating the importance of each energy component and verifying the reliability of every data point. Such work can feel exhausting and thankless, lacking the allure of novelty, yet it is undeniably essential for progress. Without these efforts, we could neither rigorously assess the capabilities of current methods nor confidently guide the development of new ones. The creation of GSCDB138 thus also reflects the collective effort of a global community devoted to accuracy and reliability in computational chemistry; without their contributions, this database would simply not have been possible.

### v2

@@ 4.3 Property-Specific Benchmark Analysis
For main-group thermochemistry (TC) (Figure 3), standard MAEs typically fall between 1 and 5 kcal/mol, indicating that even the best hybrid functionals still fail to reach chemical accuracy (1 kcal/mol) for most sets. Particularly large errors are observed for AE11 (atomic energies of heavy elements), MX34 (atomization energies of NaCl-like ionic clusters), and MB16-43 (mindless benchmarks). Ionization potential (IP) data sets, including DIPCS9 (double IP), generally show MAEs around 2–3.5 kcal/mol, while electron affinity (EA) sets average below2 kcal/mol. Among all tested functionals, CF22D and B97M-V perform best, with CF22D showing clear advantages on challenging sets like MB16-43 and TAE_W4-17MR. Notably, although B97M-V and B97X-V perform similarly on IP and EA, B97X-V outperforms B97M-V on time-dependent DFT (TDDFT) excitation energies 211, suggesting little correlation between accuracy in adiabatic IP/EA and accuracy for TDDFT excitation energies (which depends on the additional adiabatic approximation). It is possible that orbital-optimized DFT (OO-DFT) excitation energies212 might show stronger correlations. Furthermore, B97M-V and B97X-V yield reasonable performance on MX34 but greatly overestimate solid-state band gaps 213, implying that MX34 may not fully capture periodic system behavior (we also note that solid-state gaps are eigenvalue differences while MX34 are total energy differences).
@@
Barrier‑height (BH) data (Figure 5) tell a more comforting story: good hybrids sit between 1 and 2 kcal mol-1. That might look like an improvement over TC, but remember that BH reference values are usually smaller in this database, so the same absolute error can hide larger percentage mistakes. B97M‑V and CF22D perform best in this category, though CF22D shows large errors on the cycloreversion subset CRBH14.
@@
For transition metal (TM) systems (Figure 5), top-performing functionals achieve MAEs around 1–3 kcal/mol for organometallic complexes, but significantly higher errors (5 kcal/mol) occur for metal-only systems such as 3d4dIPSS, CUAGAU83, and TMD10. Performance varies with reaction type. CF22D performs poorly on reaction energies in MOR13 and ROST61, despite having the lowest overall NER for TM. In contrast, B97M-V excels for organometallic reaction energies and barriers but struggles with small cluster sets like CUAGAU83 and DAPD. Interestingly, the performance gap between empirical and non-empirical functionals narrows for TM systems, suggesting reduced advantages of empirical parameterization, at least when extrapolating beyond training data.
For Isomerization (ISO) sets (Figure 5), standard MAEs typically fall around 1–2 kcal/mol except for C20C246. B97M-V fails badly on C60ISO7 though it performs very well on other sets, which makes CF22D the best overall. consistent with known problems for range-separated hybrids on. As noted in the source study 214, 215, fullerene isomer energies are particularly challenging for functionals with a high percentage of long-range exact exchange. Interestingly, B97M(2) remedies this issue, suggesting that a balanced cancellation between exchange and correlation at medium-to-long ranges is essential. It is also worth noting that EIE22 remains challenging for semi-local functionals.
@@
For noncovalent interactions (NC) and intramolecular NC (INC) (Figure 4), sub-kcal/mol accuracy is generally attainable except for radical systems (TA13, O24) and IDISP. B97M-V, B97X-V, and B97M-V lead their respective rungs, echoing findings from MGCDB84 and GMTKN55. These functionals exhibit low errors for ionic hydration (e.g., FmH2O10, HW6Cl, HW6F), possibly reflecting overfitting. Conversely, they perform poorly on the 3B-69 set, indicating that VV10 lacks the explicit three-body terms captured by DFT-D4. When such three-body terms from B97M-D4 are included, the NERs of B97M-V improve significantly—from 1.78 to 1.21 for 3B-69 and from 1.11 to 0.98 for 3BHET. Moreover, the mean NER for INC decreases from 1.07 to 0.81, although the mean NER for NC remains largely unchanged. We therefore recommend adding three‑body dispersion to VV10‑based DFAs for large systems.
@@
For vibrational frequencies (Figure 5), the typical MAE is around 30 , with the best-performing hybrid and double hybrid functionals, PBE0‑D4 and revDSD-PBEP86-D4, reducing this to 22 and 10 , respectively. For dipole moments, the mean absolute regularized error is about 3.5%. Static polarizabilities show a standard MARE of 2.4% on Pol130 but only about 1.3% on HR46 and T144, reflecting the greater challenge of spin-polarized or exotic molecules in Pol130. For orientation-dependent electric fields (OEEF), the standard MARE rises to 8.8%, indicating that stronger fields amplify functional errors. Interestingly, for electric field properties (EF) (Figure 5), B97M-V and CF22D clearly underperform other hybrids, contrary to trends in ground-state energy performance. This reversal reinforces the concern that B97M-V and CF22D may overfit energy-based training sets and highlights the need to include molecular response properties during functional development to ensure broader transferability.
@@ 4.4 Additional Considerations for Double Hybrid Functionals
Compared to hybrid functionals, double-hybrid (DH) functionals offer substantial improvements across most data sets. The overall mean NERs of B97M(2) and revDSD-PBEP86-D4 are 0.79 and 1.52, respectively—representing relative improvements of 27% and 33% compared to their hybrid analogues, B97M-V and PBE0-D4. However, these gains come with new challenges due to the inclusion of MP2 correlation, which is a quite limited approach to wavefunction correlation. The presence of MP2 may introduces sensitivity to computational choices of certain functionals, such as basis set size, frozen-core treatment, and spin contamination (a proxy for strong correlation in many cases216, 217, 56).
@@
One important consideration is basis-set incompleteness. The MP2 component converges only as with the maximum angular momentum , whereas the SCF part approaches the CBS limit exponentially. 218 A quadruple-zeta basis is therefore still far from complete for a double hybrid. In practice one should use the same basis that was employed during parametrization. Accordingly, we pair B97M(2) with def2-QZVPPD whenever possible, since that is how the functional is defined. An instructive exception is the AE11 atomic set: a specially designed core-correlated basis reduces the error much more than def2-QZVPPD, indicating that explicit CBS extrapolation or custom basis optimization can be essential for the most demanding cases. The result also suggests that future double hybrids should be trained with larger, explicitly CBS-compatible basis sets. Two promising strategies have recently been proposed to mitigate the slow basis-set convergence of double hybrids: F12 correlation treatments 219 and density-based basis-set corrections 220, 221. Incorporating such approaches may enable the development of next-generation double hybrids with improved basis-set efficiency.
One more important source of error in DH-DFT the use of frozen core (FC) approximations. Since the semi-local component of a DH is fitted with a specific FC protocol, calculations should, in principle, replicate that protocol. For example, B97M(2) was trained with the FC approximation. If this is omitted, the MAE on the AlkAtom19 set rises dramatically from 0.18 kcal/mol to 6.24 kcal/mol—an increase by a factor of 35. However, the problem becomes more complicated when considering elements outside the training set or software-specific definitions of frozen orbitals. For instance, in Q-Chem, the K+ ion is treated with 9 frozen orbitals (a very large core!) by default, leading to an 8.15 kcal/mol MAE on the CHB6 set. Without FC, the error drops to just 1.23 kcal/mol. Another extreme case is AE11, where enabling FC inflates the error by more than two orders of magnitude. To avoid such artifacts, we report all-electron results for B97M(2) whenever the original FC recipe is demonstrably inadequate (e.g. HEAVY28_13, HEAVY28_14, CHB6_3, CHB6_6, MB16-43, AE11, MX34, and all TM-related sets).
revDSD-PBEP86-D4 was originally developed using a distinct frozen-core (FC) protocol that is difficult to automate within our current workflow. Therefore, no FC approximation is applied for revDSD-PBEP86-D4 in this study, except for the AE11 and AE18 data sets. This treatment illustrates the potential pitfalls of a non-black-box approach that requires careful handling of the FC approximation. For reference, the WTMAD2 of revDSD-PBEP86-D4 is only about 25% larger than that of B97M(2) for the thermochemistry category in GMTKN55, whereas its mean NER on TC in the present work is roughly twice as large as B97M(2). The primary source of this discrepancy is spin-symmetry breaking (discussed below), although the FC protocol also contributes significantly. Moreover, since the MP2 component in modern double-hybrid functionals—particularly RI-MP2 or localized RI-MP2—is often faster than the SCF stage, the FC approximation may not be essential for practical speedups and might best be avoided altogether in both training and production calculations. (We note, however, that this would require basis sets capable of describing core-valence and core correlation effects.)
The most critical challenge for double-hybrid density functionals (DHDFTs) is spin contamination (SC), or in other words, spin-symmetry breaking (SSB), as has already been observed in some other studies 38, 222, 56. A clear example of this issue arises for the 3B-69 data set, where revDSD-PBEP86-D4 shows abnormally large errors. The problem originates from the p-benzoquinone trimer: while the monomer is free of SC, both the dimer and trimer exhibit significant contamination, resulting in an error of about 20 kcal/mol, compared to less than 0.1 kcal/mol for PBE0-D4. To quantify its impact, we divided the 8,377 reactions into three groups according to the presence or absence of SC in r2SCAN0-D4 and revDSD-PBEP86-D4, and compared the “win rate” of revDSD-PBEP86-D4 relative to PBE0-D4 (i.e., the fraction of reactions where it yields a smaller error). For the 5,447 reactions free of SC in both functionals, the win rate is an impressive 78%. This rate drops noticeably to 64% for reactions where both functionals exhibit SC. The most severe degradation occurs for reactions in which only revDSD-PBEP86-D4 suffers from SC, likely caused by artificial symmetry breaking due to its high fraction of Hartree–Fock exchange (HFX); here, the win rate falls to just 38%, indicating that this double hybrid can underperform its hybrid counterpart in such cases.
This trend is consistent with dataset-level behavior: revDSD-PBEP86-D4 clearly outperforms PBE0-D4 on single-reference sets such as TAE_W4-17nonMR and BH28 but deteriorates significantly on strongly multireference systems, including TAE_W4-17MR and ORBH36. Its poor performance on electric-field response properties also stems from this issue. In contrast, B97M(2) is less susceptible because it uses B97M-V orbitals, suggesting that xDH-type functionals whose orbitals are generated using a lower HFX content can substantially mitigate artificial SSB. Nonetheless, the improvement of B97M(2) over B97M-V on multireference-dominated data remains smaller than for single-reference cases. Emerging approaches such as the random phase approximation (RPA) 223 or regularized MP2 methods 224, 225 behave better than MP2 in the presence of strong correlation, but cannot correct SSB in the orbitals. There is incentive to revisit the idea of orbital optimization (OO) in the presence of wavefunction correlation (the OO-DH approach226) or to include singles as well to restore broken orbital symmetries.
Our results illustrate both the current utility and present-day limitations of DH functionals. To ensure size-consistent results (because use of restricted orbitals is catastrophic for multireference problems or in heterolytic dissociations), one must use the unrestricted orbitals whenever they are lower in energy than restricted orbitals at the SCF level. This is what we have done here, with the benefit of guaranteed size-consistency, but with the drawback of reduced DH accuracy when either artificial or essential SSB occurs. Alternatively, in the regime of artificial SSB,55, 56 much better DH results could be obtained by using the (unstable) restricted orbitals. This is a user’s choice that comes at the expense of the DH no longer obeying model chemistry precepts 227: potential energy curves may not be continuous, and results may not be size-consistent. While this is disturbing, it is just like the use of standard MP2 or even RPA or CCD, whose HF orbitals are highly susceptible to artificial SSB.55 This is a manifestation of the famous symmetry dilemma of quantum chemistry.228 To minimize this problem, xDH methods with orbitals that do not suffer from extensive artificial SSB (such as B97M(2)) are preferable, as this is the primary origin of its performance advantage over revDSD-PBEP86-D4.
@@ 5 Conclusions
We have presented GSCDB137, a rigorously curated suite of 137 benchmark sets that roughly doubles the data-point count of GMTKN55/MGCDB84 (to 8377 energies) while broadening the chemical and physical scope. New components include realistic transition-metal thermochemistry and kinetics, three-body non-covalent interactions, dipoles, polarizabilities, oriented-field response energies, and dimer vibrational frequencies. At the same time, legacy data have been re-evaluated, obsolete reference values replaced, and spin-contaminated or redundant entries removed. By gathering all of these elements under a single, fully documented umbrella, GSCDB137 furnishes a coherent, high-accuracy platform for functional assessment today and a robust training ground for the next generation of non-empirical or machine-learned density functionals.
Benchmarking 29 broadly used density functionals on this database confirms the expected Jacob’s-ladder hierarchy but also highlights important exceptions. Double hybrids remain the most accurate rung, with B97M(2) and revDSD-PBEP86-D4 reducing the average error by roughly 30% relative to their hybrid analogues. Among hybrids, B97M-V emerges as the best HMGGA overall, while B97X-V is the most balanced HGGA, especially for electric-field and vibrational properties. B3LYP only performs well for frequency calculations when used with at least a triple-zeta basis set; the commonly employed B3LYP/6-31G* combination remains inadequate, as shown in previous studies.170 B97M-V is the clear leader among meta-GGAs, and revPBE-D4 is the safest choice within the conventional GGA class. However, some categories violate the ladder ordering: semi-local r2SCAN-D4 outperforms many hybrids on frequencies, and EF errors show little correlation with other ground-state energetics. These findings warn against relying on energy-only training sets and underscore the need to include density-sensitive and response quantities in future functional design. From an applications standpoint, the present study suggests the following guidelines. For main-group chemistry or organometallic reactions, B97M-V (or its D4 variant, or the runner-up, CF22D, where VV10 is unavailable) offers the best accuracy. CF22D is also a valuable choice for metal clusters. For frequencies, B3LYP-D4, PBE0-D4, and r2SCAN-D4 are recommended.
Double-hybrid accuracy is impressive but not turnkey. We show that frozen-core protocols, basis-set completeness, and especially spin symmetry breaking character can each change double-hybrid errors by a large ratio (even orders of magnitude). In practical calculations we therefore recommend (i) reproducing the training frozen-core definition whenever possible, (ii) using at least the training basis—def2-QZVPPD for B97M(2)—and (iii) using a good hybrid functional to generate orbitals to avoid artificial symmetry breaking. When such precautions are unfeasible, a robust hybrid (e.g., B97X-V or B97M-V) may be the safer alternative at present.
@@
This work was supported by the Director, Office of Science, Office of Basic Energy Sciences, of the U.S. Department of Energy through the Gas Phase Chemical Physics Program, under Contract No. DE-AC02-05CH11231. We thank the authors of MGCDB84, GMTKN55, and ACCDB for providing a solid foundation for this work. We are deeply grateful to Diptarka Hait for insightful discussions and valuable suggestions on refining the dipole moment and static polarizability data sets. We also thank Bun Chan for his generous guidance on the use of his benchmark sets and for providing so much highly accurate reference data. Our appreciation extends to Tarek Scheele and Tim Neudecker for clarifying details of the OEEF set. We also thank Klaas Giesbertz and Sebastian Ehlert for their suggestion on our preprint version of the database.
Most importantly, we thank all researchers who devote their time and effort to generating high-quality reference values, whether or not their work is included in this database. We fully recognize not only the substantial computational resources to complete high-level wavefunction calculations, as well as the human time required in the meticulous and often tedious process of evaluating the importance of each energy component and verifying the reliability of every data point. Such work can feel exhausting and thankless, lacking the allure of novelty, yet it is undeniably essential for progress. Without these efforts, we could neither rigorously assess the capabilities of current methods nor confidently guide the development of new ones. The creation of GSCDB137 thus also reflects the collective effort of a global community devoted to accuracy and reliability in computational chemistry; without their contributions, this database would simply not have been possible.